OpenStudy (anonymous):

if 29 more than 2 times a number is -15 what is the number?

OpenStudy (nerdgirl6):

This might be better suited to a math concept but I will help you. If x represents the unknown number, then you can solve the problem using the equation, 2x+29=-15 Try solving that!

OpenStudy (anonymous):

So 2x+29=-15 2x=29-15 2x=14 x=7

OpenStudy (nerdgirl6):

Careful! To take the 29 from both sides you have to subtract!